A tennis ball is dropped from 1.15 m above the
ground. It rebounds to a height of 0.993 m.
With what velocity does it hit the ground?
The acceleration of gravity is 9.8 m/s(Let
down be negative.)
1
Expert's answer
2019-09-27T09:42:53-0400

The law of conservation of energy gives


"E_i=E_f"

or

"mgh=\frac{mv^2}{2}"

So, the velocity of a ball when it hit the ground


"v=\sqrt{2gh}=\sqrt{2\times 9.8\:\rm m/s^2\times 1.15\: m}=4.75\:\rm m/s"